About 1h30m (about the time the trees should have been opened) after traffic was swung back to the p2p in scl1 loss increased on the circuit.
This bug will track this issue.
Abovenet is dispatching a tech.
As described in bug 681701, the Above folks saw some optical loss. Yesterday (10/16/2011), during the scl1 outage (bug 694877), ravi cleaned the optics and re-seated all of the connectors. We haven't seen packet loss on the interface since the changeover to the P2P ~ 17:30 yesterday.
If this continues working through tomorrow, then we can call this fixed.
The Tuesday-night failure described in bug 694877 *may* be another instance of this failure, although that is not certain. At any rate, the recommended fixes for this problem (optics replacement) are being handled in that bug.
*** This bug has been marked as a duplicate of bug 694877 ***